OpenStudy (anonymous):

solve the system be using a table -0.2x-0.3y=0.2 0.4x-0.3y=1.4

OpenStudy (anonymous):

solve first one for either x or y then put into the other equation then solve it

OpenStudy (anonymous):

sove for y it look easier 0.2x-3y=0.2 -3y=0.2-0.2x noe put this eq. into the other one 0.4x+(.2-.02x)=1.4 0.2x=1.4-.2=1.2 x=1.2/.2=6